What is the OH concentration of a 0.08M solution of CH3COONa. [Ka(CH3COOH)=1.8×10^5)

1
Expert's answer
2021-11-19T08:57:02-0500

"pH=7+ \\frac{1}{2}\n\n pK \na\n\u200b\t\n + \n\\frac{1}{2}\n\u200b\t\n logC"


"pH=7+ \\frac{1}{2}\n\u200b\t\n (5\u2212log(1.8))+ \\frac{1}{2}\n\u200b\t\n log(0.08)"


"pOH=14\u2212(7+ \\frac{1}{2}\n\u200b\t\n (5\u2212log(1.8))+ \n\\frac{1}{2}\n\n\u200b\t\n log(0.08))=5\u2212log(0.669)"


"[OH ^\n\u2212\n ]=0.669\u00d710 ^{\n\u22125}\n M"